OpenStudy (anonymous):

Please help b= 4a-17-2b

OpenStudy (anonymous):

what do you need to find?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

in terms of b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so we want to get everything around the a out the way: b= 4a-17-2b get rid of the 2b, 3b = 4a -17 get rid of the 17, 3b+17=4a now divide both sides by 4 (4b+17)/4 = a

OpenStudy (anonymous):

b = 4a - 17 - 2b b + 17 + 2b = 4a 3b + 17 = 4a (3b + 17)/4 = a

OpenStudy (anonymous):

Thank you guys so much

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

*(3b+17)/4 = a